Three uniform spheres are located at the corners of an equilateral triangle. Each side of the triangle has a length of 2.80 m. Two of the spheres have a mass of 2.10 kg each. The third sphere (mass unknown) is released from rest. Considering only the gravitational forces that the spheres exert on each other, what is the magnitude of the initial acceleration of the third sphere?
_______m/s2
Answer #1

Force, F1 = G m1 m2/r^2

F1 = (6.67 x 10^-11 x 2.1 x m) /2.8^2 = 1.79 x 10^-11 m

Force, F2 = F1 = 1.79 x 10^-11 m

Force, F = sqrt (F1^2 + F2^2 + 2 F1 F2 cos60)

F = 3.09 x 10^-11 m N

Acceleration, a = F/m = 3.09 x 10^-11 m/s^2

Comment in case any doubt please rate my answer.....